• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Carlos

    Log entry time 08:18:32 on October25,2004

    Entry number 135201

    keyword=Owl shift summary

    Shift crew:
    C. Munoz Camacho
    C. Ferdi

    Shift summary:

    00:00 Shift started with beam in the hall running production in kinematic#3.

    00:32 MCC is taken beam away in order to send pulsed beam to HallB, according to the plan.

    01:15 Beam restored to the hall.

    Quite and peaceful night.

    07:00 Beam taken away for maintenance.
    Calorimeter LED run.

    08:15 Started cosmics run.

    08:20 Move target to empty.